The acceptable tolerance on a shaft product is 9 ± 0.005 in. From past experience, it is known that σ= 0.003 in. What is the percentage of scrap parts?
Answer #1

Solution:-

P(scarp park) = P(X < 8.995 or X > 9.005)

= P(X < 8.995) + P(X > 9.005)


Z for 8.995 = (8.995-9)/0.003 = -1.667

P value = 0.047757

=> P(X < 8.995) = 0.047757 = 0.0478

Z for 9.005 = (9.005-9)/0.003 = 1.667


P value = 0.9522


P(X < 9.005) = 0.9522


P(X > 9.005) = 1-0.922 = 0.0478


=> P(Score park) = P(X < 8.995 or X > 9.005)
= P(X < 8.995) + P(X > 9.005)
= 0.0478 + 0.0478
= 0.0956